# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

### v2

@@ 4.2 Computations
As a first step of the analysis, the overall temporal trend, $\widehat{X}_{it}$, is obtained as the fitted values from the model in Equation (4), where $h({\bf s}_{i},t;df)$ represents a principal spline of space and time (see, e.g., Mardia et al., 1998, Sahu and Mardia, 2005, Fontanella et al., 2013) with $df$ bases. This two-step strategy should be preferred to a joint modeling of the exposure and the outcome, as it allows avoiding feedback from the outcome to the exposure that may lead to biased estimators of the exposure effect (see, e.g., Stephens et al., 2023).
An MCMC scheme is then implemented to simulate from the posterior distribution of $\bm{\Theta}$. The algorithm proposed by Chan and Jeliazkov (2009) is used to sample from the full conditionals of the time-varying coefficients. It is an efficient simulation algorithm that allows to sample all the regression coefficients (for all the time instants) in a single step. The remaining full conditional distributions are inverse-gamma distributions. The latent process $\vartheta_{it}$ is an exception and is sampled using Metropolis-Hastings steps. To avoid identifiability issues, all time-specific, area-specific and interaction effects are centered at zero during each MCMC iteration (see, e.g., Knorr-Held, 2000, Franco-Villoria et al., 2022). Further details can be found in Section C of the Supplementary Material.
@@ 5 Competing Approaches
In this section, we briefly describe the competing approaches that are considered in the simulation study (in Section 6) and in the real data application (in Section 7). Similarly to SDGLMC, they are all based on the assumption that the outcome follows a Poisson distribution. The main difference lies in how the exposure effect is modeled, and how unmeasured confounding is accounted for.
The linear predictor is a function of exposure and measured confounders. To control for unmeasured confounding, additional covariates and basis functions are included in the linear predictor (see, e.g., Peng et al., 2006, Peng and Dominici, 2008). Several approaches from the literature are then fitted under the Bayesian paradigm: (i) the Null model, which does not account for unmeasured confounding; (ii) the GLMadj model, which adjusts for unmeasured confounding through a natural cubic spline of time (see, e.g., Peng and Dominici, 2008, Zhou et al., 2021, Masselot et al., 2022); (iii) the JDZ model by Janes et al. (2007), which is similar to the previous approach but further decomposes the exposure effect in two orthogonal parts; (iv) the Dummy model (see, e.g, Dai et al., 2014, Klompmaker et al., 2023), which allows for season-specific exposure effects; and (v) the Periodic model by Peng et al. (2005) and Chen et al. (2013), which assumes that the exposure effect is a linear combination of trigonometric functions. A hierarchical structure is assumed for each of them, with spatial-specific coefficient vectors being independently and identically distributed. Further details are provided in Section D of the Supplementary Material.
@@ 6 Analysis of Artificial Data
To validate the proposed approach, we conduct two simulation studies, both inspired by the real data example presented in Section 2. In the first one, which is described next, we want to assess the ability of SDGLMC to mitigate the confounding bias.
@@
We consider the ${N}=117$ contiguous HDs of Piemonte and Lombardia, then we assume that ${T}=600$ and there are no measured confounders (i.e., ${p_{m}}=0$). The exposure $X_{it}$ and the unmeasured confounder $Z_{it}$ are correlated vector autoregressive processes of order $1$ (VAR($1$)) with a linear correlation coefficient of 0.5, and with spatially-structured, temporally independent innovations (Cressie and Wikle, 2015, Sahu, 2022). A proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022) is assumed to model the spatial dependence of these innovations. The marginal spatio-temporal means of exposure and unmeasured confounder are generated from Fourier basis expansions. The correlation matrix of $X_{it}$ is parameterized by spatial and temporal autocorrelation parameters, i.e. $\phi_{x}^{S}\in(0,1)$ and $\phi_{x}^{T}\in(0,1)$, respectively. Similarly, $\phi_{z}^{S}$ and $\phi_{z}^{T}$ represent the autocorrelation parameters of $Z_{it}$, which is sampled conditional on $X_{it}$. These four parameters represent the strength of the spatial and temporal dependence, and define three scenarios that are summarized in Table 1.
It is worth noting that, in the concurvity setting, $X_{it}$ and $Z_{it}$ are considered as fixed unknown functions, particularly when exposure and unmeasured confounder vary at large scales. The stochastic generative model we use continues to be valuable within this framework of fixed functions because the realizations of $X_{it}$ and $Z_{it}$ provide reasonable values for exposure and unmeasured confounder that could arise in real-world applications where a reasonable stochastic process is absent (Paciorek, 2010).
@@
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,,}$
@@
The proposed approach is compared to the Null model and to the GLMadj model. The results from the other competing approaches are not reported here, but they can be found in Section E of the Supplementary Material. A natural cubic spline of time with $6$ degrees of freedom is used to account for unmeasured confounding in GLMadj. Following Peng et al. (2006), the degrees of freedom for the spline are chosen by fitting several models with different numbers of bases and selecting the minimum number from which the estimates appear stabilized.
Figure 3 depicts the results in terms of percent change of risk associated with a 10-unit increase in the exposure, specifically focusing on its constant and temporal components, i.e. it is defined as $P_{t}=100(\exp\{10(\overline{\delta}_{1}+\delta_{1t})\}-1)$. For each scenario, the black solid line represents the space-averaged true exposure effect. The estimate of $P_{t}$ obtained from SDGLMC is shown in red, with the solid line indicating its average, and the shaded area indicating the percentiles $2.5$ and $97.5$ of its distribution over 100 replicates.
@@
The dashed gray and dashed green lines represent the pooled estimates from the Null and GLMadj models, averaged over the 100 replicates. Whereas the Null model is evidently biased in all cases, GLMadj and SDGLMC significantly reduce the confounding bias in Scenario S1, although without entirely removing it. In Scenario S2, the confounding bias reduction is still achieved by SDGLMC compared to the Null model, but to a lesser extent. This could be because the unmeasured confounder varies at high frequencies in time, so it cannot be captured by the time-varying intercepts that only account for gradual changes. Also, the difference $X_{it}-\widehat{X}_{it}$ is approximately as rough as $Z_{it}$, and, due to their correlation, a portion of the variability of the unmeasured confounder is explained by the local exposure effect. Scenario S3 is very similar to Scenario S1, suggesting that the lack of smoothness in space of the unmeasured confounder does not seem to alter the performance (in terms of confounding bias) of the proposed approach.
@@
Figure 3: Simulation Study 1: space-averaged percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$, for three scenarios and for the Null (dashed gray line), GLMadj (dashed green line), and SDGLMC (red solid line) models. The shaded area represents the percentiles $2.5$ and $97.5$ of the posterior means of $P_{t}$ obtained from SDGLMC over 100 replicated datasets. The black solid line represents the true exposure effect.
With regard to the variability of the exposure effect over space, we have computed the posterior means of the spatial effects, $\tilde{\bm{\delta}}_{1}$, for each scenario. The average of these posterior means over 100 replicated datasets are shown in Figures 46 and are compared to the (averaged) posterior means of the differences between area-specific and pooled estimates – i.e., the same quantity depicted in Figure 9 – obtained from the competing models.
@@
All methods recover the simulated spatial pattern, with lower values in the middle of the study area. SDGLMC provides accurate estimates in all scenarios, while the Null and GLMadj models tend to underestimate the spatial variability in some cases.
Figure 4: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S1.
Figure 5: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S2.
Figure 6: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S3.
@@
The results above are confirmed also by Tables E.4–E.6 in the Supplementary Material, where the average mean squared error ($MSE_{avg}$), the maximum absolute bias ($MAB$), and the average mean absolute error ($MAE_{avg}$) are reported for each model and each scenario (see the Supplementary Material, Section E.2, for definitions). We also report these metrics for each component of the exposure effect to assess their contribution to the overall error. In general, we note that SDGLMC outperforms the competing models in all scenarios.
@@
In the second simulation study, described in Section F of the Supplementary Material, we want to assess whether the proposed approach is able to recover the first derivative, $f^{\prime}(\cdot;\bm{\gamma}_{f})$, when the functional form is assumed known. The data are generated using Equations (1)–(2), and we show that SDGLMC is able to capture the dynamics of the exposure effect, but there remains a residual bias which is due to confounding.
Another objective of this second simulation study is to motivate, from a practical viewpoint, the decomposition of the exposure effect into local-scale and large-scale components described in Equation (3). We compare SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. As expected from the discussion in Section 3, the results suggest that the proposed approach is more effective in reducing the confounding bias and always has lower $MSE_{avg}$, $MAB$ and $MAE_{avg}$ values.
@@ 7 Application to the Italian Mortality Data
We apply SDGLMC to the Italian mortality dataset (see Section 2), and compare its results to those from the competing approaches. The large-scale trend, $\widehat{X}_{it}$, is obtained using a principal spline with $5$ spatial bases and $20$ temporal bases. For each HD and time instant, ${\bf M}_{it}$ contains bases from a natural cubic spline of daily temperature and daily relative humidity (with 7 and 4 degrees of freedom, respectively), and indicator variables for weekdays, holidays and lockdown periods (hence ${p_{m}}=19$; see also Section A of the Supplementary Material).
@@
We can envision several distributional specifications for the space-time interaction component of the intercept, $\delta_{0it}^{*}$, depending on the correlation structure we want to impose. Knorr-Held (2000) proposes four types of models, giving rise to four types of space-time interaction terms, characterized by different prior assumptions. These are reviewed by Franco-Villoria et al. (2022) and Orozco-Acosta et al. (2023), and are summarized in Table 2. Type T1 assumes unstructured space-time variability, and is obtained by setting $\delta_{0it}^{*}=0$ for all $i$ and $t$, since an unstructured error is already present in Equation (7). Type T2 assumes independent temporal trends for each HD, with a common variance for the innovations. Type T3 assumes independent spatial trends for each time instant, with a common conditional variance. Type T4 assumes complete dependence, i.e. the temporal trends are different from HD to HD, but are more likely to be similar if the HDs are neighbors.
The prior assumptions discussed in Knorr-Held (2000) can be further relaxed. In fact, during the pandemic, some HDs in Lombardy – more specifically in the province of Bergamo – experienced a much greater increase in deaths than any other HD. Therefore, a fifth interaction type, T5, is also considered, assuming independent space-time interactions across the HDs that follow random walk processes with area-specific evolution variances. This captures temporal trends that differ considerably among HDs but do not have any spatial structure.
Table 2: Space-time interaction types considered for $\delta_{0it}^{*}$. DIC values for the corresponding models are reported, with the lowest value highlighted in italics.
@@
After fitting a model for each of the five types of space-time interaction components, the one with the lowest deviance information criterion (DIC; Spiegelhalter et al., 2002) is selected. Based on the last column of Table 2, the model with interaction type T5 has the smallest DIC value, hence its results will be discussed below. A reason for this may be that, in comparison to the models proposed by Knorr-Held (2000), the model with T5 seems more able to capture very specific spatio-temporal trends, such as those observed in the Bergamo area due to the pandemic.
@@
We assess the adequacy of the model fit by analyzing Bayesian $p$-values for each HD and time instant. The chosen measure evaluates how well the model aligns with the observed data, and involves calculating the quantity $\Pr\bigl{(}Y_{it}>Y^{rep}_{it}\bigr{)}+\frac{1}{2}\Pr\bigl{(}Y_{it}=Y^{rep}_{it}\bigr{)}$, where $Y^{rep}_{it}$ is generated from the predictive distribution in each MCMC iteration. Bayesian $p$-values close to $0.5$ suggest that the generated data align well with the model. Conversely, values approaching the boundaries of the unit interval, i.e. $<0.1$ or $>0.9$, indicate a poor fit (Gelman et al., 2013). The estimated mean $p$-value is around $0.5$ suggesting an adequate fit.
@@
Figure 7 depicts the estimated space-averaged local exposure effect (i.e., $\overline{\delta}_{1}+\delta_{1t}$), expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red solid line represents its posterior mean, while the shaded area the corresponding 95% credible interval (CI). The exposure effect is not constant in time, as no horizontal straight line can be drawn within the CI. This effect, instead, exhibits varying peaks across different years, and three phases are clearly identifiable: before (2018-19), during (2020-21) and after (2022) the pandemic outbreak.
@@
During the first two years, the local exposure effect exhibits a seasonal pattern, with the peaks occurring during the summer months. The estimated association is bounded away from zero from August to September 2018 from April to October 2019, with the highest change in mortality risk of $5.50\%$ (95% CI: $[3.03\%,7.98\%]$) occurring on June 29, 2019.
Figure 7: Time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations estimated by SDGLMC. The red solid line represents its posterior mean, while the shaded area the corresponding 95% CI.
It is worth noting that a higher percent change of mortality risk occurs when the fine particulate matter is generally less present in the air, so this may indicate that a threshold below which the exposure effect is null might not exist (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). The seasonal behavior seems to be confirmed by other studies (Nawrot et al., 2007, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018).
These seasonal changes of the risk suggest that the unmeasured confounders could have not only a marginal effect on mortality (hence influencing $\beta_{0it}$), but also an interaction effect with the exposure (hence affecting $\beta_{1it}$). In fact, the seasonal pattern might be due to potential interactions between PM2.5 concentrations and air temperature. As noted in Section 1, increased ventilation and physiological susceptibility to air pollution may explain the higher percent change of risk during summer. Besides, interaction with other variables, such as other pollutants, are also possible (see, e.g., Liu et al., 2023, Ji et al., 2022, Siddika et al., 2019).
Previous studies have also reported that particle composition may modify the effect of PM2.5 on mortality. In fact, fine particulate matter is a complex mixture of organic (such as black carbon and organic carbon) and inorganic compounds (such as sulfate, nitrate, and ammonium), and the relative proportion of these components may vary in time (Franklin et al., 2008, Masselot et al., 2022, Vinson et al., 2020). Indeed, their precursors are emitted from both natural (e.g., oceans and volcanoes) and anthropogenic sources, such as fossil fuel combustion, industrial processes, and residential heating, and their transportation and transformation to PM2.5 in the atmosphere are mainly influenced by meteorological conditions (Hao et al., 2023, Jing et al., 2020, Kelly and Fussell, 2012). Toxicological studies indicate that the PM constituents have different effects on human health (see, e.g., Kelly and Fussell, 2012), and some authors have suggested that ammonium is probably among the most harmful components (Masselot et al., 2022, Hao et al., 2023).
During the pandemic phase, the CI always contains the zero except for few days during December 2021. The literature reveals a lack of consensus regarding the association between air pollution and adverse COVID-19 health outcomes in the general population (see, for example, Brunekreef et al., 2021a, Walton et al., 2021). More recently, the review by Bhaskar et al. (2023) suggests a positive association in the majority of previous studies. Our findings, instead, indicate that the local exposure effect in 2020 and 2021 is generally not bounded away from zero (except for few days). To justify this, consider first of all that our model uses all-cause mortality in the elderly population as the health outcome, so it is not directly comparable to the COVID-19 studies mentioned earlier. Secondly, the issuance of the lockdown measures by the Italian government, as well as the mandatory use of facial masks, not only effectively curbed the transmission of the SARS-CoV-2 virus (and of other viruses), but also potentially reduced exposure to air pollution. During the lockdown periods, people were required to stay at home, resulting in reduced exposure to outdoor air pollutants. The ventilation was limited because the lockdowns occurred mainly during winter months. Moreover, the use of facial masks may have reduced the inhalation of airborne hazardous particles (Kodros et al., 2021, Zhou et al., 2018).
Although lockdown periods were not issued in the second half of 2021, the local exposure effect remains not bounded away from zero. This could be due to the fact that people continued to adopt measures such as wearing masks and sanitizing hands. Additionally, it is possible that the phenomenon of mortality displacement or harvesting may have occurred. This is a consequence of the fact that the most vulnerable people may have died earlier than expected in the preceding years due to the SARS-CoV-2 virus in 2020-2021, resulting in an apparently null (or even protective) effect of the exposure afterwards (Armstrong et al., 2017, Bhaskaran et al., 2013, Zanobetti and Schwartz, 2008, Zanobetti et al., 2000).
During the third phase, people started to return to their normal activities, resulting in increased exposure to air pollution. In fact, the local exposure effect has a shape similar to that in the first phase, and is bounded away from zero from July to September 2022, with the highest change in mortality risk of $3.53\%$ (95% CI: $[0.61\%,6.44\%]$) occurring on July 26.

### v3

@@ 4.2 Computations
The inference procedure starts by first estimating the overall temporal trend $\widehat{X}_{it}$ as the fitted values from Equation (4), where $h({\bf s}_{i},t;df)$ represents a principal spline of space and time (e.g., Mardia et al., 1998, Sahu and Mardia, 2005, Fontanella et al., 2013) with $df$ bases. This strategy is preferred over a joint modeling of the exposure and the outcome because it avoids feedback from the outcome to the exposure that may lead to biased estimators of the exposure effect (e.g., Stephens et al., 2023).
A Markov chain Monte Carlo (MCMC) scheme is implemented to simulate from the posterior distribution of $\bm{\Theta}$. The algorithm proposed by Chan and Jeliazkov (2009) is used to sample from the full conditionals of the spacetime-varying coefficients. It is an efficient simulation algorithm that allows to sample the coefficients for all instants of time in a single step. To avoid identifiability issues, all time-specific, area-specific and interaction effects are centered at zero during each MCMC iteration (e.g., Knorr-Held, 2000, Franco-Villoria et al., 2022). The latent process $\vartheta_{it}$ is sampled using Metropolis-Hastings steps. The remaining full conditional distributions are inverse-gamma distributions. Further details about the MCMC algorithm can be found in Section C of the Supplementary Material.
@@ 5 Competing Approaches
This section briefly describes the competing approaches considered in the simulation study (in Section 6) and in the real data application (in Section 7). Similarly to SDGLMC, they are all based on the assumption that the outcome follows a Poisson distribution. The main differences lie in how the exposure effect is modeled and how unmeasured confounding is handled.
In each approach, the linear predictor is a function of exposure and measured confounders. To control for unmeasured confounding, additional covariates and basis functions are included in the linear predictor (e.g., Peng et al., 2006, Peng and Dominici, 2008). The following models are fitted under the Bayesian paradigm: (i) the Null model, which does not account for unmeasured confounding; (ii) the GLMadj model, which adjusts for unmeasured confounding through a natural cubic spline of time (e.g., Peng and Dominici, 2008, Zhou et al., 2021, Masselot et al., 2022); (iii) the JDZ model by Janes et al. (2007), which is similar to (ii) but further decomposes the exposure effect into two orthogonal parts; (iv) the Dummy model (e.g., Dai et al., 2014, Klompmaker et al., 2023), which allows for season-specific exposure effects; and (v) the Periodic model by Peng et al. (2005) and Chen et al. (2013), which assumes that the exposure effect is a linear combination of trigonometric functions. Each model adopts a hierarchical structure, where spatial-specific coefficient vectors are assumed to be independently and identically distributed. Further details are provided in Section D of the Supplementary Material.
@@ 6 Analysis of Artificial Data
We conduct two simulation studies, inspired by the real data example presented in Section 2, to validate the proposed approach.
@@
The first study assesses the ability of SDGLMC to mitigate confounding bias. We consider the N=117{N}=117 contiguous HDs of Piemonte and Lombardia. We assume T=600{T}=600 and ${p_{m}}=0$ (i.e., there are no measured confounders). The exposure, $X_{it}$, and the unmeasured confounder, $Z_{it}$, are generated as correlated vector autoregressive processes of order 11 [VAR(11)] with a linear correlation coefficient of $0.5$. For both processes, the innovations are temporally independent but spatially structured. The spatial dependence is parameterized using a proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022). The marginal spatio-temporal means of the exposure and the unmeasured confounder are generated from Fourier basis expansions. The correlation structure of $X_{it}$ is parameterized by spatial and temporal autocorrelation parameters, i.e., $\phi_{x}^{S}\in(0,1)$ and $\phi_{x}^{T}\in(0,1)$, representing the strength of the spatial and temporal dependence, respectively. Similarly, $\phi_{z}^{S}$ and $\phi_{z}^{T}$ represent the autocorrelation parameters of $Z_{it}$. Then, $Z_{it}$ is sampled conditionally on $X_{it}$. The parameters $\phi_{x}^{S}$, $\phi_{x}^{T}$, $\phi_{z}^{S}$, and $\phi_{z}^{T}$ are varied to define three scenarios that are summarized in Table 1.
@@
$\vartheta_{it}=o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,$,
@@
SDGLMC is compared to the Null model and to the GLMadj model here. The results from the other competing approaches can be found in Section E of the Supplementary Material. A natural cubic spline of time with $6$ degrees of freedom is used to account for unmeasured confounding in GLMadj. This choice follows from fitting several models with different numbers of bases and selecting the smallest number from which the estimates appear stabilized (Peng et al., 2006).
The results are reported in Table 2 and Figures 3 and 4 and are expressed as percent change in risk associated with a 10-unit increase in the exposure. For each scenario, the reported values summarize the distribution of the posterior means over the 100 replicates, namely their average and the $2.5$ and $97.5$ percentiles are taken into consideration.
@@
Table 2 compares the pooled estimates from the Null and GLMadj models to the baseline effect estimated from SDGLMC. Considering that the simulated value is $5\%$, the Null model exhibits a significant positive bias in all scenarios. GLMadj and SDGLMC effectively reduce (although without entirely removing) confounding bias in Scenario S1. In Scenario S2, the bias reduction is less pronounced. A possible explanation is that the unmeasured confounder varies at high frequencies in time, making it difficult to be adequately captured by time-varying intercepts (or natural splines of time in the case of GLMadj) that only account for gradual changes. The difference $X_{it}-\widehat{X}_{it}$ is approximately as rough as $Z_{it}$ in time. Also, as all these variables are correlated, part of the variability of the unmeasured confounder is explained by the local exposure effect. Performance under Scenario S3 is very similar to Scenario S1, suggesting that the lack of smoothness in space of the unmeasured confounder does not seem to alter the performance in terms of confounding bias of the proposed approach.
Table 2: Simulation Study 1: baseline component of the exposure effect under each scenario, expressed as percent change in risk associated with a 10-unit increase in the exposure. Values represent the average posterior means across 100 replicates, with $2.5$ and $97.5$ percentiles in brackets. The true simulated effect is $5\%$.
Scenario Null GLMadj SDGLMC
S1 10.427 [9.507, 11.310] 5.189 [4.233, 5.872] 5.282 [5.020, 5.568]
S2 8.394 [8.118, 8.713] 7.061 [6.809, 7.528] 7.145 [6.849, 7.637]
S3 10.028 [9.406, 10.688] 5.462 [5.120, 6.286] 5.568 [5.217, 6.320]
Figure 3 depicts the temporal component of the exposure effect. For each scenario, the black solid line represents the true effect. The estimate obtained from SDGLMC is shown in red, with the solid line indicating its average over 100 replicates, and the shaded area indicating the percentiles $2.5$ and $97.5$. The proposed approach successfully recovers the simulated temporal pattern in all scenarios, as opposed to the Null and GLMadj models, which do not account for time-varying effects.
@@
Figure 3: Simulation Study 1: temporal component of the exposure effect, expressed as percent change in risk associated with a 10-unit increase in the exposure, for each scenario. The red solid line represents the estimate (averaged over 100 replicates) obtained from SDGLMC, and the shaded area indicates the percentiles $2.5$ and $97.5$. The black solid line represents the true temporal component. Null and GLMadj models are not shown because they assume an exposure effect that is constant in time.
@@
Figure 4 shows the spatial component of the exposure effect for each scenario. The (averaged) posterior means obtained from SDGLMC are compared to the (averaged) posterior means of the differences between area-specific and pooled estimates obtained from the competing approaches. All methods generally recover the simulated spatial pattern, with lower values in the middle of the study area. SDGLMC provides accurate estimates in all scenarios, while the Null and GLMadj models tend to underestimate the spatial effect in some districts.
Figure 4: Simulation Study 1: spatial component of the exposure effect, expressed as percent change in risk associated with a 10-unit increase in the exposure. The top panel shows the true spatial component. The remaining panels show the posterior means (averaged over 100 replicates) under each scenario (rows) and model (columns).
@@
The above results are also confirmed by Tables E.4–E.6 in the Supplementary Material, where the average mean squared error ($MSE_{avg}$), the maximum absolute bias ($MAB$), and the average mean absolute error ($MAE_{avg}$) are reported for each model and each scenario (see Section E.2 in the Supplementary Material for definitions). We also report these metrics for each component of the exposure effect to assess their contribution to the overall error. In general, we note that SDGLMC outperforms the competing models in all scenarios.
@@
The second simulation study is fully described in Section F of the Supplementary Material. A first objective is to assess whether the proposed approach is able to recover the first derivative $f^{\prime}(\cdot;\bm{\gamma}_{f})$ when the data are generated using Equations (1)–(2). We show that SDGLMC is able to capture the dynamics of the exposure effect, although a residual bias remains due to confounding.
Another objective of the second simulation study is to provide practical motivation for the decomposition of the exposure effect into local-scale and large-scale components in Equation (3). We compare SDGLMC with a reparameterized model that does not remove the large-scale spatio-temporal trend from the exposure. As expected from the discussion in Section 3, the results suggest that the proposed approach is more effective in reducing the confounding bias and always has lower $MSE_{avg}$, $MAB$ and $MAE_{avg}$ values (see Table F.7 in the Supplementary Material).
@@ 7 Application to the Italian Mortality Data
We apply SDGLMC to the Italian mortality dataset (see Section 2) and compare its results to those obtained from the competing approaches. The large-scale trend, $\widehat{X}_{it}$, is estimated using a principal spline with $5$ spatial basis dimensions and $20$ temporal basis dimensions. For each HD and instant of time, ${\bf M}_{it}$ includes bases from natural cubic splines of daily temperature and daily relative humidity (with 7 and 4 degrees of freedom, respectively), along with indicator variables for weekdays, holidays and lockdown periods (thus, ${p_{m}}=19$; see also Section A of the Supplementary Material).
@@
We can consider several distributional specifications for the space-time interaction component of the intercept, $\delta_{0it}^{*}$, depending on the correlation structure we want to impose. Knorr-Held (2000) proposes four types of models, each defining a different space-time interaction term characterized by distinct prior assumptions. These models are reviewed by Franco-Villoria et al. (2022) and Orozco-Acosta et al. (2023) and summarized in Table 3. Type T1 assumes unstructured space-time variability and is obtained by setting $\delta_{0it}^{*}=0$ for all $i$ and $t$ as an unstructured error is already present in Equation (6). Type T2 assumes independent temporal trends for each HD, with a common variance for the innovations. Type T3 assumes independent spatial trends for each time $t$, with a common conditional variance. Type T4 assumes complete dependence, i.e., the temporal trends are different across HDs but are more likely to be similar for neighboring HDs.
The prior assumptions discussed in Knorr-Held (2000) can be further relaxed. In fact, during the pandemic, some HDs in Lombardia — more specifically in the province of Bergamo — experienced a much greater increase in deaths than any other HD. To capture such heterogeneity, we introduce a fifth interaction type, T5, which assumes independent space-time interactions across HDs following random walk processes with area-specific evolution variances. Thus, interaction T5 captures temporal trends that differ considerably among HDs but do not have any spatial structure.
Table 3: Space-time interaction types considered for $\delta_{0it}^{*}$. DIC values are reported for model selection; the lowest value is in italics.
@@
After fitting a model for each of the five types of space-time interaction components, the one with the lowest deviance information criterion (DIC; Spiegelhalter et al., 2002) is selected. As shown in the last column of Table 3, the model with interaction type T5 has the smallest DIC value, probably because it is able to capture highly specific spatio-temporal trends—such as those observed in the Bergamo area during the pandemic. Thus, the results of the model with interaction type T5 are discussed in detail.
@@
To assess the fit of the posterior distribution of SDGLMC, we compare the observed data to the posterior predictive distribution. For each HD and time point, we compute a Bayesian $p$-value based on the test quantity $\Pr\bigl(Y_{it}>Y^{rep}_{it}\bigr)+\frac{1}{2}\Pr\bigl(Y_{it}=Y^{rep}_{it}\bigr)$, where $Y^{rep}_{it}$ is generated from the posterior predictive distribution at each MCMC iteration. Values near $0.5$ suggest that the model provides an adequate fit to the observed data. Conversely, values approaching the extremes of the unit interval (i.e., <0.1<0.1 or >0.9>0.9) indicate poor fit (Gelman et al., 2013). The estimated mean $p$-value is approximately $0.5$, supporting the adequacy of the model.
@@
Figure 5 depicts the estimated space-averaged local exposure effect, expressed as the percent change in risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. In other words, the quantity $100(\exp\{10(\overline{\delta}_{1}+\delta_{1t})\}-1)$ is computed for each MCMC iteration, which represents the average effect across all HDs at a given time $t$. The red solid line represents the posterior mean, and the shaded area denotes the 95% posterior credible interval (CI). The exposure effect is not constant in time as no horizontal straight line can be drawn within the CI. Instead, it exhibits distinct peaks across different years, and three phases can be identified: 2018-19, 2020-21 and 2022.
Figure 5: Posterior summary (solid line: mean, shaded area: point-wise limits of 95% posterior credible interval) of the space-averaged local exposure effect, expressed as percent change in mortality risk for a $10\ \mu g/m^{3}$ increase in PM2.5 concentrations. Vertical dashed lines mark calendar years.
@@
During the first two years, the local exposure effect exhibits a seasonal pattern, with the peaks occurring during the summer months. The estimated association is bounded away from zero between August and September 2018 and from April to October 2019, with the highest change in mortality risk of $5.50\%$ (95% CI: $3.03\%,7.98\%]$) observed on June 29, 2019. In the second phase (2020-21), the CI generally includes zero, except for a few days toward the end of the period. Finally, in the third phase (2022), the shape of the local exposure effect resembles that of the first phase, remaining bounded away from zero between July and September, with the highest change in mortality risk of $3.53\%$ (95% CI: $0.61\%,6.44\%]$) occurring on July 26.